Please help me solve to select the set of equations that represents the following situation: The cost of six CDs and four DVDs is $77.80; and the cost of 14CDs and 11 CDs is $192.25. How much do a CD and a DVD cost?
6x-4y=77.80; 14x^11y=192.25
6x+4y=77.80; 14x+11y=192.25
<6x^4y=77.80; 11x^14y=192.25
6x+4y=77.80; 11x+14y=192.25

Answer by jorel1380(3719)

6x+4y=77.80
14x+11y=192.25
--------------
66x+44y=855.80
56x+44y=769
10x=86.8
x=8.68
y=6.43
CD's cost $8.68;DVD's cost $6.43..